OpenStudy (anonymous):

A motorboat that travels with a speed of 20km/hr is in still water has traveled 36 km against the current and 22km with the current, having spent 3 hours on the entire trip. Find the speed of the current of the lake.

OpenStudy (anonymous):

@jim_thompson5910 @ganeshie8 @Mertsj Help

jimthompson5910 (jim_thompson5910):

Let b = speed of boat in still water c = speed of current

jimthompson5910 (jim_thompson5910):

since "A motorboat that travels with a speed of 20km/hr is in still water", we know that b = 20

jimthompson5910 (jim_thompson5910):

now set up the distance/rate/time equations for upstream and downstream and solve for t in each case upstream: d = rt 36 = (b - c)*t 36 = (20 - c)*t t = 36/(20-c) downstream: d = rt 22 = (b + c)*t 22 = (20 + c)*t t = 22/(20+c)

jimthompson5910 (jim_thompson5910):

now the two times must add to 3 hours, so this means time1 + time2 = 3 36/(20-c) + 22/(20+c) = 3 solve for c